Circular No. 8512

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bk IN MCG +07-33-27 R. A. Jansen and K. Tamura, Arizona State University; and N. A.

Grogin, Johns Hopkins University, report their discovery of an apparent supernova in UBVRI CCD images obtained with the 1.8-m Vatican Advanced Technology Telescope atop Mt. Graham. Jansen first found the conspicuous point-like object at R = 17.95 ± 0.05 on two 300-s R-band images taken on Apr. 2.46 UT, it and was confirmed on images taken by Jansen and Tamura on Apr. 4.4, 5.4, 6.4, and 7.5. The new object is located at α = 16^{h}02^{m}17^{s}.04, δ = +42^{o}54′55′′.3 (equinox 2000.0), which is 4′′.5 east and 5′′.6 south of the center of the nucleus of MCG +07-33-27. Nothing is visible at this location on a Sloan Digital Sky Survey image or on an R-band CCD image obtained by Grogin on 1995 May 31.4 with the F. L. Whipple Observatory 1.2-m telescope (limiting mag R ≃ 21.3). There is a nearby bright (red mag ∼ 11, blue mag ∼ 13) foreground star with position end figures 09^{s}.94, 04′′.6.

SUPERNOVA 2005bl IN NGC 4059 Further to IAUC 8507, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Apr. 14.34 (at mag 18.8) and 15.36 UT (mag 18.3). SN 2005bl is located at α = 12^{h}04^{m}12^{s}.26, δ = +20^{o}24′24′′.8 (equinox 2000.0), which is 13′′.2 east and 11′′.4 south of the nucleus of NGC 4059. A KAIT image taken on Mar. 11.33 showed nothing at this position (limiting mag 19.5).

COMET 9P/TEMPEL J. Crovisier, P. Colom, N. Biver, D. Bockelée-Morvan, and A. Leca-

cheux, Observatoire de Paris, report: “The OH lines at 18 cm were observed in comet 9P/Tempel (the target of the ‘Deep Impact’ mission) with the Nançay radio telescope. The average spectrum from Mar. 20 to Apr. 14 shows detection of the 1667-MHz line with an area of 12 ± 2 mJy km/s, corresponding to an OH production rate of 4 × 27 molecules/s.”

Visual total-magnitude estimates: Mar. 8.57 UT, 12.7 (S. Yoshida, Ibaraki, Japan, 0.40-m reflector); 11.83, 12.9 (A. Pearce, Noble Falls, W. Australia, 0.20-m reflector); 20.18, 12.6 (N. Biver, Poigny-la-foret, France, 0.26-m reflector); 29.87, 11.8 (A. Baransky, Pylypovychi, Ukraine, 0.36-m reflector); Apr. 4.57, 11.5 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 10.86, 11.4 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 12.14, 11.6 (J. J. Gonzalez, León, Spain, 0.20-m reflector).

2005 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green